# EchoTorrent: Towards Swift, Sustained, and Streaming Multi-Modal Video Generation

### v1

@@ 4.1 Experimental Settings
Implementation. The teacher and student models share this baseline architecture from InfiniteTalk [81]. For Streaming Enhancement Initialization, we initialize the student model from LightX2V, an autoregressive video diffusion model. The generated videos have a spatial resolution of $400\times 720$ and a total duration of 3 latent frames. For ACC-DMD, we set the audio CFG scale to 1.8. In autoregressive streaming, each generated chunk consists of 4 latent frames, and we maintain a KV cache of the most recent 12 frames. To balance causality and quality, we convert causal attention layers to bidirectional attention in a staggered pattern (every other layer), resulting in a hybrid attention mechanism. Experiments are conducted on 64 PPUE GPUs, with 5K iterations for Streaming Enhancement Initialization, 10K iterations for ACC-DMD, 5K iterations for Hybrid Long Tail Forcing, and 2K iterations for Extra-latent Distillation, respectively. The batch size is set to 1 for each GPU. The learning rate is set to 2e-6 for the student and 4e-7 for the fake score.
@@ 5.1 Multimodal Video Generation
Video Generation. The field of video generation has evolved from early 2D U-Net architectures augmented with handcrafted motion priors to modern DiT diffusion frameworks [22, 67, 23, 83, 88, 43, 29, 24, 49, 7, 90, 31, 44, 38, 52, 80, 77, 77, 2, 1]. While spatiotemporal convolutions enabled models such as SVD [3] to generate short clips with high spatial fidelity, transformer-based diffusion models have significantly advanced temporal coherence and motion smoothness by replacing hard-coded inductive biases with data-driven self-attention mechanisms. Recent progress in text-to-video generation is largely driven by large-scale pretraining [51, 62, 53, 93, 46, 70, 69, 42, 64, 74]. CogVideoX [83] achieves fine-grained semantic alignment through hierarchical token fusion; HunyuanVideo [32] introduces a dual-stream DiT architecture to enhance conditional control robustness; and Wan [65] demonstrates that scaling model capacity alone can yield state-of-the-art visual fidelity, setting new benchmarks for general-purpose video synthesis.
Audio-Driven Avatar Animation. Audio-driven avatar generation aims to synthesize expressive and temporally coherent human performances from speech signals [19, 40, 48, 63, 28, 12, 59, 71, 57, 27, 35, 36, 17, 89, 66, 33, 94, 47, 37]. While early approaches primarily focused on talking heads [95, 78], recent advances have extended to full-body animation with improved realism and control. Representative methods include EMO [61], which employs temporal encoding for dynamic expression modeling; AniPortrait [73], which leverages 3D-to-2D pose guidance for natural body motion; Vlogger [96], which adopts diffusion-based rendering for high-fidelity synthesis; and EchoMimic [10], which enables multimodal conditioning for diverse driving inputs. CyberHost [39] further supports rich user controls such as hand gestures and body trajectories, while OmniHuman [40] improves long-term stability via progressive dropout. For multi-character scenarios, MultiTalk [58] and HunyuanVideo-Avatar [9] introduce interaction-aware modeling, and OmniAvatar [18] proposes efficient audio injection into transformer blocks for low-latency inference.
@@ 6 Conclusion
In this work, we propose EchoTorrent, a 4-NEF streaming framework for high-quality, long-horizon, multi-modal video generation. We introduce a teacher-student-Decoder collaborative post-training paradigm, design a novel hybrid causal-bidirectional architecture, and propose two novel distillation approaches (i.e., ACC-DMD and Long Tail Forcing) to address degradation under joint constraints of few NEFs, streaming mode, and long-sequences. Comprehensive experiments and ablation studies demonstrate that EchoTorrent exceeds current state-of-the-art methods in both quantitative metrics and qualitative evaluations, achieving superior performance in real-time, high-quality, and long-horizon multi-modal condition human video generation.

### v3

@@ 4.1 Experimental Settings
Implementation. The teacher and student models share this baseline architecture from InfiniteTalk [82]. After Multi-Teacher Training, we adapt the teacher and student models from LightX2V  [^1] https://huggingface.co/Kijai/WanVideo_comfy/tree/main/Lightx2v, a few-step image-to-video diffusion model. The generated videos have a spatial resolution of $400\times 720$ and a total duration of 3 latent frames. For ACC-DMD, we set the audio CFG scale to 1.8. In autoregressive streaming, each generated chunk consists of 4 latent frames, and we maintain a KV cache of the most recent 12 frames. To balance causality and quality, we convert causal attention layers to bidirectional attention in a staggered pattern (every other layer), resulting in a hybrid attention mechanism. Experiments are conducted on 64 PPUE GPUs, with 5K iterations for teachers training, 10K iterations for ACC-DMD, 5K iterations for Hybrid Long Tail Forcing, and 2K iterations for VAE Decoder refining, respectively. The batch size is set to 1 for each GPU. The learning rate is set to 2e-6 for the student and 4e-7 for the fake score.
@@ 5.1 Multimodal Video Generation
Video Generation. The field of video generation has evolved from early 2D CNN architectures  augmented with handcrafted motion priors to recent DiT diffusion frameworks [22, 68, 23, 84, 89, 43, 29, 24, 49, 7, 91, 31, 44, 38, 53, 81, 78, 78, 2, 1]. While spatiotemporal convolutions enabled models such as SVD [3] to generate short clips with high spatial fidelity, transformer-based diffusion models have significantly advanced temporal coherence and motion smoothness by replacing hard-coded inductive biases with data-driven self-attention mechanisms. Recent progress in text-to-video generation is largely driven by large-scale pretraining [52, 63, 54, 94, 46, 71, 70, 42, 65, 75]. CogVideoX [84] achieves fine-grained semantic alignment through hierarchical token fusion; HunyuanVideo [32] introduces a dual-stream DiT architecture to enhance conditional control robustness; and Wan [66] demonstrates that scaling model capacity alone can yield state-of-the-art visual fidelity, setting new benchmarks for general-purpose video synthesis.
Audio-Driven Avatar Animation. Audio-driven avatar generation aims to synthesize expressive and temporally coherent human performances from speech signals [19, 40, 48, 64, 28, 12, 60, 72, 58, 27, 35, 36, 17, 90, 67, 33, 95, 47, 37]. While early approaches primarily focused on talking heads [96, 79], recent advances have extended to full-body animation with improved realism and control. Representative methods include EMO [62], which employs temporal encoding for dynamic expression modeling; AniPortrait [74], which leverages 3D-to-2D pose guidance for natural body motion; Vlogger [97], which adopts diffusion-based rendering for high-fidelity synthesis; EchoMimic [10] which enables multimodal conditioning for diverse driving inputs; and EchoMimicV2 [51], which first extends face generation to upper-body human generation. CyberHost [39] further supports rich user controls such as hand gestures and body trajectories, while OmniHuman [40] improves long-term stability via progressive dropout. For multi-character scenarios, MultiTalk [59] and HunyuanVideo-Avatar [9] introduce interaction-aware modeling, and OmniAvatar [18] proposes efficient audio injection into transformer blocks for low-latency inference.
@@ 6 Conclusion
In this work, we propose EchoTorrent, a 4-NFE streaming framework for high-quality, long-horizon, multi-modal video generation. We introduce a teacher-student-Decoder collaborative post-training paradigm, design a novel hybrid causal-bidirectional architecture, and propose two novel distillation approaches (i.e., ACC-DMD and Long Tail Forcing) to address degradation under joint constraints of few NFE, streaming mode, and long-sequences. Comprehensive experiments and ablation studies demonstrate that EchoTorrent exceeds current state-of-the-art methods in both quantitative metrics and qualitative evaluations, achieving superior performance in real-time, high-quality, and long-horizon multi-modal human video generation.